If 2x/3 =18 what is the value of 2x - 5
Answer by jim_thompson5910(35256)

First isolate x in 2x/3 = 18

2x/3 = 18
2x = 18*3
2x = 54
x = 54/2
x = 27

Now use this to compute 2x-5

2x-5 = 2(27)-5 = 54-5 = 49

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------